​(Common stock valuation​) Abercrombie​ & Fitch's common stock pays a dividend of $1.25. It is currently selling for $34.53 . If the​ firm's investors require a return of 8 percent on their investment from buying Abercrombie​ & Fitch​ stock, what growth rate would Abercrombie​ & Fitch have to provide the​ investors?

The growth rate Abercrombie​ & Fitch would have to provide the investors is ?
Answer #1

Last Dividend, D0 = $1.25
Current Price, P0 = $34.53
Required Return, rs = 8.00%

Let growth rate be g%

D1 = D0 * (1 + g)
D1 = $1.25 * (1 + g)

rs = D1 / P0 + g
0.0800 = $1.25 * (1 + g) / $34.53 + g
0.0800 = 0.0362 * (1 + g) + g
0.0800 = 0.0362 + g * 0.0362 + g
0.0438 = 1.0362 * g
g = 0.0423 or 4.23%

Growth Rate = 4.23%

The growth rate Abercrombie​ & Fitch would have to provide the investors is 4.23%